the sum of three consecutive integers is 1035. find the integers.
Answer by Alan3354(30993)

the sum of three consecutive integers is 1035. find the integers.
------------
1035/3 = 345, the middle integer
--------------
You can do:
n + n+1 + n+2 = 1035 and find that n = 344
--------------
You can do:
(n-1) + n + (n+1) = 3n = 1035
--> n = 345
==============
Pick one